OpenStudy (anonymous):

In triangle ABC, what is the measure of A if a = 11, b = 13, and c = 7? 57.79° 49.62° 38.17° 32.21°

OpenStudy (campbell_st):

use the law of cosines... \[\cos(A) = \frac{c^2 + b^2 - a^2}{2bc}\]

OpenStudy (anonymous):

ight

OpenStudy (anonymous):

7^2+13^2-11^2/2(13)(7)?

OpenStudy (campbell_st):

that will give the cos(A) to find A \[A = \cos^{-1}(answer)\]

OpenStudy (anonymous):

so 182

OpenStudy (anonymous):

97/182?

OpenStudy (campbell_st):

no... use brackets (b^2 + c^2-a^2)/(2bc)

OpenStudy (campbell_st):

ok... then find the answer

OpenStudy (campbell_st):

oops angle

OpenStudy (anonymous):

I multiplied the bottom and got 182

OpenStudy (campbell_st):

ok... so on the calculator is shift cos then 97/182

OpenStudy (anonymous):

and then I get 53.3%

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

well that's not correct... does you calculator have a small D at the top so you are working in degrees..?

OpenStudy (anonymous):

nope

OpenStudy (campbell_st):

what does it have,,?

OpenStudy (anonymous):

when I do (cos)^-1 97/182 I get 90

OpenStudy (anonymous):

90 degrees

OpenStudy (campbell_st):

here is an online calculator http://web2.0calc.com/ press these keys 2nd acos 2nd 97 divide 182 ) then =

OpenStudy (anonymous):

and then I get 58.75 degrees

OpenStudy (anonymous):

Its 57

OpenStudy (anonymous):

57.79